If 250 kJ of energy are absorbed in the following reaction, how many moles of CO2 reacted:

6CO2 + 6H2O → C6H12O6 + 6O2 ΔH=2803 kJ



1
Expert's answer
2020-10-05T14:26:07-0400

Let's make a proportion according to the reaction equation:

n = 6 moles of CO2 — ΔH=2803 kJ;

n1= x moles of CO2 — ΔH1=250 kJ.


"n1 = \frac{n*ΔH1}{ΔH};
n1 = \frac{6*250}{2803} = 0.5351 kJ."